OpenStudy (anonymous):

For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 4x - y = 8 2x + 3y = 9 Answer -8 + 4x 8 + 4x -8 - 4x 8 - 4x

OpenStudy (anonymous):

-8+4x

OpenStudy (anonymous):

to isolate y in 4x - y = 8 we get y by itsself, so we move the 4x term over to the other side by subtracting it from both sides, now we have -y = -4x + 8 divide or multiply by negative 1 to get rid of the negative sign and get y = 4x - 8 now plug that into the other one 2x + 3y = 9 like this 2x + 3(4x - 8) = 9 Distribute 2x + 12x - 24 = 9 14x = 33 x = 33/14